Greg's family rode in a taxicab last satuday night. The total cost of the ride, with a 20% tip, was $25.80. What was the cost of the ride without tip?
Answer by bucky(2189)

Let C represent the cost of the ride without a tip. So when Greg's family arrived at their destination, the meter said that they owed the driver C.
.
They apparently got good service so they decided to add a 20% tip. That means that they multiplied the cost on the meter by 0.2 (0.2 is equivalent to 20%). and they added that to C. When they did that the total was $25.80. In equation form you can write that:
.
C + 0.2*C = $25.80
.
You can add the two terms on the left side because they both involve C. The 1*C plus 0.2*C equals 1.2*C and this makes the equation become:
.
1.2*C = $25.80
.
Now you can solve for C by dividing both sides of the equation by 1.2 to get:
.
C = $25.80/1.2 = $21.50
.
So you now know that C (the cost of the ride) was $21.50. To that amount they added a 20% tip (0.2 times $21.50 = $4.30). So the $21.50 for the ride and the $4.30 for the tip add up to $21.50 + $4.30 = $25.80, just as the problem said it should.
.
In summary, the cost of the taxicab ride was $21.50
.
Hope this helps you to understand the problem a little better.
.